OpenStudy (anonymous):

What is the area of a square with a diagonal of the square root of 2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):



OpenStudy (anonymous):

a. 1/2 b. 1 c. \[\sqrt{2}\] d. 2 e. 4

hero (hero):

area = 1x1 = 1

OpenStudy (anonymous):

diagonal of the square to a side is \[\sqrt{2}\] so if the diagonal is the square root of 2, the side is 1 and the square has area 1

OpenStudy (anonymous):

(V2) ^2 = 2a^2 -> a^2 = 1 Thus A = a2 = 1

Directrix (directrix):

Also, the area of a square is half the product of its 2 congruent diagonals. A = (1/2) √2 * √2 = (1/2) *2 = 1